Solve for x.

12^(sqrt{x^2}) - 24^(x - 2) = 144

Answer by ikleyn(52790)

.
Solve for x.

12^(sqrt{x^2}) - 24^(x - 2) = 144
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

It is not for algebraic solution - in Algebra, there are no methods solving such equations.


Use either your graphics calculator or online tools to get approximate numerical solutions.


For example, you may use an online solver https://www.desmos.com/calculator


Print there  y = 12^(sqrt{x^2}) - 24^(x - 2)  in one window and  y = 144  in another window.



Calculator will provide the plot, and you can identify two root  

    x = -2 (an approximate value)  and x = 2.00281  (another approximate value) 


by clicking at the intersection points.   ANSWER

Solved.